### v1

# Rewarding Graph Reasoning Process makes LLMs more Generalized Reasoners

Abstract.
Despite significant advancements in Large Language Models (LLMs), developing advanced reasoning capabilities in LLMs remains a key challenge. Process Reward Models (PRMs) have demonstrated exceptional promise in enhancing reasoning by providing step-wise feedback, particularly in the context of mathematical reasoning. However, their application to broader reasoning domains remains understudied, largely due to the high costs associated with manually creating step-level supervision. In this work, we explore the potential of PRMs in graph reasoning problems - a domain that demands sophisticated multi-step reasoning and offers opportunities for automated step-level data generation using established graph algorithms. We introduce GraphSilo, the largest dataset for graph reasoning problems with fine-grained step-wise labels, built using automated Task-oriented Trajectories and Monte Carlo Tree Search (MCTS) to generate detailed reasoning steps with step-wise labels. Building upon this dataset, we train GraphPRM, the first PRM designed for graph reasoning problems, and evaluate its effectiveness in two key settings: inference-time scaling and reinforcement learning via Direct Preference Optimization (DPO). Experimental results show that GraphPRM significantly improves LLM performance across 13 graph reasoning tasks, delivering a 9% gain for Qwen2.5-7B and demonstrating transferability to new graph reasoning datasets and new reasoning domains like mathematical problem-solving. Notably, GraphPRM enhances LLM performance on GSM8K and MATH500, underscoring the cross-domain applicability of graph-based reasoning rewards. Our findings highlight the potential of PRMs in advancing reasoning across diverse domains, paving the way for more versatile and effective LLMs. Models and data are available at https://huggingface.co/GraphPRM.
Large Language Models, Graph Reasoning, Process Reward Model
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Natural language processing
1. Introduction
Despite remarkable progress in scaling Large Language Models (LLMs) and their performance on established benchmarks, developing advanced reasoning capabilities—particularly in domains such as mathematical problems and code generation—remains a critical area of research (Cobbe et al., 2021a; Wang et al., 2023b; Zhou et al., 2023; Li et al., 2023; Zheng et al., 2023; Gou et al., 2024; Chen et al., 2023; Liu et al., 2025b). Chain-of-Thought (CoT) (Wei et al., 2022; Yao et al., 2023) generation plays a critical role in facilitating such reasoning by breaking down complex problems into intermediary steps, thus emulating human cognitive processes. Recent research has highlighted the success of prompting LLMs with CoT examples and finetuning them using question-CoT solution pairs (Yuan et al., 2023; Chen et al., 2024a; Yuan et al., 2024; Li et al., 2025; Liu et al., 2025a), with the latter method proving particularly effective.
Currently, with the release of O1/3-like models, there has been increasing attention on the viability of using reward models to provide supervisory signals within reinforcement learning (Christiano et al., 2017; Ouyang et al., 2022; Rafailov et al., 2023; Lightman et al., 2024; Jiao et al., 2024), as a means to enhance the reasoning capabilities of LLMs, especially in optimizing test-time computing (Wang et al., 2023c; Snell et al., 2024; Kumar et al., 2024). Two main categories of reward models have been identified: Outcome Reward Models (ORMs) (Cobbe et al., 2021b; Yu et al., 2024), which deliver feedback solely at the conclusion of the problem-solving process, and Process Reward Models (PRMs) (Li et al., 2023; Uesato et al., 2022; Lightman et al., 2024; Wang et al., 2024b), which provide detailed feedback at every step of reasoning. The ORM evaluates the entire sequence to assign a score, while the PRM assesses each step of the reasoning process individually. PRMs have several unique advantages. They give detailed feedback by identifying where errors occur, which is crucial for reinforcement learning and correcting mistakes automatically. Moreover, PRMs mimic human evaluation processes, understanding that an intermediate error leads to a faulty conclusion, similar to human judgment. To date, PRMs have shown exceptional performance on complex reasoning by offering fine-grained supervision.
However, the exploration of PRMs has largely been confined to mathematical reasoning tasks (Wang et al., 2024b; Wu et al., 2023), as shown in Figure 1. This narrow focus limits the understanding of how reward-based enhancements can be applied across diverse reasoning domains. Moreover, training PRMs to help mathematical problem-solving often involves constructing highly structured step-level data, which is challenging to automate due to the significant manual effort requirement (Lightman et al., 2024; Zheng et al., 2024). This scalability constraint, combined with the domain-specific nature of current PRM applications, raises important questions about their broader applicability and practical implementation across different reasoning contexts.
Figure 1. Example of LLM reasoning with process supervision on mathematical problems and graph reasoning problems.
Graph computational problems (GCPs), by contrast, offer a promising avenue to overcome this limitation. GCP reasoning (Chen et al., 2024a; Tang et al., 2024b; Li et al., 2024c; Zhang et al., 2024; Li et al., 2024b; Peng et al., 2024; Xu et al., 2023) involves tackling challenges related to graph theory, a branch of mathematics concerned with the study of graph structures made up of nodes (vertices) connected by edges. GCP encompasses a wide range of complex tasks, including pathfinding, network analysis, and edge counts, which require sophisticated multi-step reasoning and the ability to navigate intricate relational structures. GCPs are intrinsically challenging for LLMs due to their dependence on spatial reasoning, the need to maintain and manipulate multiple interrelated nodes, and the requirement for precise logical deductions. Furthermore, many traditional graph algorithms already have well-established solutions. These can be leveraged to automatically generate step-level data, offering a unique opportunity to build training data for LLMs that bypasses the need for manual annotation. This enables us to explore automated solution paths for GCPs while utilizing reward models to refine the reasoning process.
In light of these insights, this study explores the application of PRMs within GCP tasks to strengthen LLM. The investigation is guided by two primary research questions: First, can we train effective Graph Process Reward Models that significantly bolster LLMs’ graph computational problem-solving capabilities? Second, do these GCP-specific reward models demonstrate versatility by improving LLM performance in other complex reasoning tasks? Addressing these questions will provide valuable insights into the scalability and cross-domain applicability of reward-based enhancements in LLMs, contributing to developing more versatile reasoning LLMs.
To this end, we first introduce the GraphSilo dataset—the foremost training dataset designed specifically for training PRMs in GCP tasks. GraphSilo represents the largest collection of GCPs with CoT solutions available, and it stands out as the sole process supervision dataset that incorporates detailed step-level correctness labels for GCPs. To build this dataset, we employ Monte Carlo Tree Search (MCTS) (Luo et al., 2024a) alongside Task-oriented Trajectories, enabling the automated generation of detailed reasoning steps (More details in Section 3.1). Building on GraphSilo, we then train the first process reward model tailored for GCP tasks, named GraphPRM. We demonstrate the efficacy of GraphPRM in two main settings: 1) Inference-Time Scaling: GraphPRM is used to rank multiple responses generated by the base LLM, selecting the best answer during inference. 2) Reinforcement Learning: GraphPRM helps to identify high-quality, step-wise reasoning solutions by distinguishing between correct and incorrect reasoning steps for GCPs. Positive-negative pairs are constructed based on these distinctions, with the correct reasoning steps serving as positive examples and the incorrect ones as negative examples. This pairing enables the creation of preference data for Direct Preference Optimization (DPO) (Rafailov et al., 2023). DPO is then applied to further refine the model’s reasoning ability by optimizing the preference for high-quality solutions over suboptimal ones.
Experimentally, with the help of GraphPRM, we observe significant improvements across three different backbone LLMs, including Qwen2.5 (Team, 2024b), Llama3.1 (AI@Meta, 2024), and Gemma 2 (Team, 2024a). For example, GraphPRM helps Qwen2.5-7B achieve a 9% improvement across 13 GCP tasks during inference-time computation. Additionally, GraphPRM demonstrates impressive transferability: 1) GCP Dataset Transfer: In zero-shot settings, it enables different LLMs to achieve a 20-30% improvement on the GraphWiz (Chen et al., 2024a) test set across various GCP tasks, and 2) Reasoning Domain Transfer: Most notably, when applied to mathematical reasoning tasks, GraphPRM boosts LLM performance on four math datasets, such as GSM8K and Math-500, enhancing their mathematical reasoning capabilities. These results confirm that rewarding graph reasoning processes improve LLMs’ reasoning abilities, advancing both graph computational problem-solving and other reasoning domains like mathematical problem-solving. Ultimately, our findings show that integrating graph-based reward models like GraphPRM makes LLMs more versatile and effective across various reasoning tasks, bridging gaps between graph computation and general reasoning applications.
Figure 2. Overall illustration of process annotation pipeline of GraphSilo (a, b) and the strategies of utilizing GraphPRM to strengthen LLMs on graph reasoning, including scaling inference-time computation (c) and reinforcement learning (d).
2. Preliminaries and Related Works
2.1. LLMs for Graph Reasoning
Recent research has focused on integrating LLMs with graph data due to their growing prominence. Early works like NLGraph (Wang et al., 2023a) and GPT4Graph (Guo et al., 2023) evaluated LLMs’ ability to understand graph-structured data. Subsequent studies explored fine-tuning LLMs for graph tasks, leading to models like GraphLLM (Chai et al., 2023) and GraphGPT (Tang et al., 2024a). Graph computational problems, requiring deeper structural understanding and multi-step reasoning, remain particularly challenging. GraphInstruct (Luo et al., 2024b) evaluated LLMs’ understanding of graphs and improved LLMs’ performance on graph reasoning tasks. GraphArena (Tang et al., 2024b) introduced a benchmark to test LLMs’ reasoning abilities with graph computational challenges, while ProGraph (Li et al., 2024a) proposed a challenging benchmark for graph analysis using external APIs. Further advancements have focused on improving graph reasoning capabilities. GraphQA (Fatemi et al., 2024) showed that graph encoding significantly impacts LLM reasoning. GraphWiz (Chen et al., 2024a) improved reasoning through instruction tuning and preference alignment. Additionally, NLGift (Zhang et al., 2024) and GraCoRe (Yuan et al., 2025) provided evaluation suites to test generalization and reasoning with graph data. Despite significant efforts, LLMs are still struggling with solving complex graph computation problems, highlighting the need for more fine-grained, step-level process supervision in graph reasoning processes.
2.2. Reward Models for Reasoning Problem
Reasoning - the capacity to derive new insights, tackle challenges, and reach conclusions from given information - stands as a core capability of LLMs. Researchers have proposed various methods to improve or elicit the reasoning ability of LLMs (e.g., mathematical reasoning), which can be broadly divided into three groups: pre-training (OpenAI, 2023; AI@Meta, 2024; Team, 2024b, a), fine-tuning (Yu et al., 2024; Luo et al., 2023; Yue et al., 2024; Chen et al., 2024c) and prompting (Imani et al., 2023; Wang et al., 2023c; Chen et al., 2024b). Except for directly improving and eliciting the reasoning potential of LLMs, researchers have been dedicated to boosting LLM reasoning results via an extra verifier for selecting the best answer from multiple decoded candidates. There are two primary types of verifiers: the Outcome Reward Model (ORM) (Cobbe et al., 2021b) and the Process Reward Model (PRM) (Lightman et al., 2024; Wang et al., 2024b; Zheng et al., 2024).
Outcome Reward Model Given a graph computational problem $p$ and its corresponding solution $s$, ORM aims to assign a single score to $s$, which ranges in $[0,1]$ and indicates the correctness of solution $s$ towards problem $p$. Specifically, ORM is trained with a cross-entropy loss (Li et al., 2023) as follows:
(1) $\mathcal{L}_{ORM}=y_{s}\log r_{s}+\left(1-y_{s}\right)\log\left(1-r_{s}\right),$
where $y_{s}$ is the ground-truth of problem $p$, and $y_{s}=1$ if solution $s$ is correct ($y_{s}=0$ if $s$ is incorrect). $r_{s}$ indicates the sigmoid score assigned by ORM on solution $s$ from LLM generator. The performance of the ORM is critically dependent on the construction of a high-quality training dataset with outcome labels.
Level Task Definition Time Complexity Answer Type
Node # Degree Return the degree of node $u$ in a given graph $\mathcal{G}$. $O(|\mathcal{V}|)$ Integer
# Clustering Coefficient Calculate the ratio of actual edges to possible edges among node $u$’s neighbors in a graph $\mathcal{G}$. $O(|\mathcal{V}|^{2})$ Float
# Neighbor Find the neighbors of node $u$ in a given graph $\mathcal{G}$. $O(|\mathcal{V}|)$ Node List
# Page Rank Identify the node with the highest PageRank value in a given graph $\mathcal{G}$. $O(|\mathcal{V}|+|\mathcal{E}|)$ Integer
# Predecessor Return the list of predecessor nodes of $u$ in a given graph $\mathcal{G}$ $O(|\mathcal{V}|)$ Node List
# Jaccard Calculate the Jaccard Coefficient value $J$ between node $u$ and node $v$, where $J=\frac{|\mathcal{N}(u)\cap\mathcal{N}(v)|}{|\mathcal{N}(u)\cup\mathcal{N}(v)|}$. $O(|\mathcal{V}|)$ Float
Node # Common Neighbor Find the number of common neighbors of node $u$ and node $v$ in a given graph $\mathcal{G}$. $O(|\mathcal{V}|)$ Integer
Pair # Connectivity Determine if there is a path from node $u$ to node $v$ in a given in a given graph $\mathcal{G}$. $O(|\mathcal{V}|+|\mathcal{E}|)$ Boolean
# Maximum Flow Calculate the maximum amount of flow from node $u$ to node $v$ in a given graph $\mathcal{G}$. $O(|\mathcal{V}|^{2}\sqrt{|\mathcal{E}|})$ Integer
Graph # Breadth First Search Return the sequence of nodes in the order of a BFS traversal starting from node $u$. $O(|\mathcal{V}|+|\mathcal{E}|)$ Node List
# Cycle Determine if there exists an cycle in a given graph $\mathcal{G}$. $O(|\mathcal{V}|+|\mathcal{E}|)$ Boolean
# Diameter Calculate the maximum distance over all pairs of nodes in a given graph $\mathcal{G}$. $O(|\mathcal{V}|(|\mathcal{V}|+|\mathcal{E}|)\text{log}|\mathcal{V}|)$ Integer
# Minimum Spanning Tree Calculate the total weight of the minimum spanning tree in a given graph $\mathcal{G}$. $O(|\mathcal{E}|\text{log}|\mathcal{E}|)$ Integer
Table 1. Descriptions of graph tasks in GraphSilo with problem definition, time complexity of representative algorithms and answer type. $|\mathcal{V}|$ and $|\mathcal{E}|$ indicate the number of nodes and edges in the graph $\mathcal{G}$.
Process Reward Model Different from ORM, PRM incorporates the concept of supervising intermediate reasoning steps in addition to the final outcome, enabling more granular feedback throughout the reasoning process. Given a problem $p$ and solution $s^{*}$ containing multiple reasoning steps, PRM is designed to assign each step of solution $s^{*}$ a score ranging from 0 to 1. PRM is trained as follows:
(2) $\mathcal{L}_{PRM}=\sum_{i=1}^{N}y_{s_{i}}\log r_{s_{i}}+\left(1-y_{s_{i}}\right)\log\left(1-r_{s_{i}}\right),$
where $s_{i}$ is the $i$-th step of $s^{*}$, and $y_{i}$ is the ground-truth label of step $s_{i}$. $r_{s_{i}}$ indicates the sigmoid score of $s_{i}$ assigned by PRM. Likewise, the training of PRM can be regarded as a binary classification task. which means the value of $y_{i}$ can be 0 or 1. To obtain high-quality process labels of PRM training dataset, existing methods include 1) human annotation (Lightman et al., 2024) and 2) automate annotation via Monte Carlo estimation (Wang et al., 2024b; Luo et al., 2024a).
2.3. Task Formulation
To evaluate the effectiveness of our proposed GraphPRM, two aspects are taken into consideration: (1) Inference-Time scaling: we utilize GraphPRM to rank and select the correct one from multiple outputs from LLMs during inference (2) Reinforcement Learning: GraphPRM is utilized to further improve LLM in graph reasoning tasks via reinforcement learning with process supervision.
Scaling Inference-Time Performance In general, leveraging additional computation at inference time allows LLMs to surpass their training performance, unlocking new possibilities in reasoning tasks (Lightman et al., 2024; Snell et al., 2024). By integrating a PRM, the performance of an LLM on a specific prompt can be enhanced during inference. Through multiple samples generated by the base LLM, PRMs evaluate and score individual steps in each solution, employing step-wise and inter-answer aggregation to identify the optimal answer. This approach effectively scales LLM inference performance via PRM-based search. We adopt the Best-of-N and Beam Search paradigms (Snell et al., 2024) to demonstrate how PRMs improve LLM performance.
Reinforcement Learning with Process Supervision An enhanced reward model is instrumental in training higher-performing LLMs. We further use GraphPRM to boost LLMs in graph reasoning via Direct Preference Optimization (DPO) with process supervision. We utilize GraphPRM to select high-quality verified step-wise reasoning solutions with positive-negative pairs for GCPs. Then, the LLM is further optimized on the preferred reasoning solutions to further sharpen its graph reasoning abilities.
3. Methodology
In this section, we present the construction pipeline of GraphSilo, including graph task selection, graph generation, and process annotation. We then describe the supervised training of GraphPRM. Using the trained PRM model, we further introduce approaches of employing GraphPRM to enhance LLMs by scaling inference-time computation and enabling reinforcement learning training with process supervision. An overall illustration is detailed in Figure 2.
3.1. GraphSilo Construction
To enhance LLM reasoning in solving GCPs, we create a new dataset, GraphSilo, for training a PRM. This section outlines the procedures including graph task selection, graph generation, process annotation, and GraphSilo statistics.
3.1.1. Graph Task Selection
We comprehensively consider a broad range of thirteen graph reasoning tasks in GraphSilo, including node level, node pair level and graph level graph computational problems. Specifically, we include five node level tasks: Degree, Clustering Coefficient, Neighbor, Page Rank and Predecessor; four node pair level tasks: Jaccard, Common Neighbor, Connectivity and Maximum Flow; and four graph level tasks: Breadth First Search, Cycle, Diameter and Minimum Spanning Tree. The tasks are summarized in Table 1. These thirteen graph tasks encompass extensive computational and mathematical reasoning, requiring LLMs to comprehend graph structures and perform reasoning. They facilitate a deeper theoretical understanding of graph algorithms while also addressing a wide range of practical applications.
3.1.2. Graph Generation
GraphSilo features diverse graphs of varying structures and sizes to evaluate LLM capabilities. Following (Luo et al., 2024b), we include three types of graph structure distributions: 1) random networks, 2) small-world networks (Grabow et al., 2015), and 3) BA scale-free networks (Barabási and Albert, 1999). For each type, we consider both directed and undirected graphs with random generation. In terms of graph size, we include four distinct levels: Tiny (5-7 nodes), Small (8-15 nodes), Medium (16-25 nodes), and Large (26-35 nodes). These sizes represent varying levels of difficulty for LLMs in understanding and reasoning about graphs. Detailed descriptions of graph structure are presented at Appendix A.1.
Figure 3. Illustration of generating task-oriented trajectories with step-wise labels based on graph algorithm codes.
3.1.3. Process Annotation
Training an effective PRM for GCP tasks relies on step-level reasoning labels. Unlike prior work (Lightman et al., 2024), which depends on human annotations, GCP offers two inherent advantages: 1) their outcomes are explicit and well-defined, and 2) they involve analyzing and manipulating graph structures, in which established graph algorithms are efficiently used. Based on these, we propose our process annotation approach as follows:
Automated labeling with Task-oriented Trajectories Since graph algorithms follow standardized pipelines, the process of solving graph problems can naturally facilitate the generation of task-specific trajectories, which in turn produce fine-grained step-wise labels. Specifically, we select representative algorithms and corresponding codes for each GCP. Then, we translate the whole execution process of the code to trajectories in natural language, which can be further divided into multiple sub-steps, serving as positive labels. To construct negative sub-steps, these positive sub-steps can then be perturbed using three strategies: 1) structure-based: introducing errors related to the graph structure (e.g., incorrect assumptions about whether the graph is directed or connected); 2) node/edge-based: introducing errors by randomly adding or removing nodes or edges; and 3) calculation-based: introducing errors through incorrect formulas or miscalculated results.
We present a more detailed example of generating reasoning process and step-level labels in diameter task in Figure 3. In this example, the trace breaks down the steps: 1) Graph Structure Check; 2) Shortest Path Calculation; 3) Distance Calculation; 4) Error Identification; 5) Diameter Calculation. We can obtain the correct steps by printing the execution process in the code (green color). Besides, we can collect negative steps by randomly introducing structure errors in graph structure check part, node/edge errors in the shortest path calculation part, and calculation errors in the diameter calculation part (red color). With obtained reasoning processes and step-wise labels, we employ Qwen2.5-72B-Instruct to rephrase them, mimicking LLM distributions while avoiding uniform templates.
Dataset Training Set CoT Process Step-wise Label Task Size
NLGraph (Wang et al., 2023a) 8 29,370
GraphQA (Fatemi et al., 2024) 12 2,300
GraphInstruct (Luo et al., 2024b) 21 18,900
GraphWiz (Chen et al., 2024a) 9 72,785
NLGift (Zhang et al., 2024) 4 37,000
ProGraph (Li et al., 2024a) 3 29,260
GraphArena (Tang et al., 2024b) 10 10,000
GraphSilo (Ours) 13 118,189
Table 2. Comparison of GraphSilo with related datasets.
Automated labeling via Monte Carlo Method To enhance the limited variety of task-oriented trajectories from graph algorithms, we use the Monte Carlo method (Wang et al., 2024b; Luo et al., 2024a) to automatically generate and label graph reasoning steps. Following (Luo et al., 2024a), we construct a search tree for each problem-answer pair, where each node contains the problem $p$, solution prefix $s$, and and all rollouts $\{(s,r_{i})\}^{k}_{i=1}$ derived from that node (where $r_{i}$ denotes the $i$-th rollout). Each edge in the tree represents either a single step or a sequence of steps originating from the node. To automatically collecting graph reasoning process with labels, we adhere to the principle of selecting the most valuable rollouts during tree search. Thus, for each node, we calculate both its value function $Q(s,r)$ and exploration term $U(s)$. The value function is defined as $Q(s,r)=\alpha^{1-\mathrm{MC}(s)}\cdot\beta^{\frac{\operatorname{len}(r)}{L}}$, where $\alpha,\beta$ and $L$ are hyperparameters, $\operatorname{len}(r)$ denotes the number of tokens in a rollout, and $\mathrm{MC}(s)$ is the Monte Carlo estimation. $Q$ is supposed to indicate how likely a rollout will be chosen for each iteration. Additionally, the exploration term is computed as: $U(s)=c_{\text{puct}}\frac{\sqrt{\sum_{i}N\left(s_{i}\right)}}{1+N(s)}$, where $N(s)$ denotes the visit count and $c_{\text{puct}}$ is a constant determining the level of exploration. During the selection phase, a rollout is popped and selected according to tree statistics based on the PUCT algorithm: $(s,r)=\arg\max_{(s,r)}[Q(s,r)+U(s)]$. A binary search is subsequently employed to locate the first error in the selected rollouts, adding rollouts with $0<MC(s)$ to the candidate pool. All positions preceding the first error are treated as new states for further exploration.
3.1.4. GraphSilo Statistics
This extensive dataset provides a strong foundation for training step-wise PRM and evaluating LLMs on various GCP tasks. As shown in Table 2, our dataset is the largest one to date containing CoT solutions, with 118,189 samples and 394,165 step-wise labels. It is also the only dataset that includes step-level supervision, offering a valuable resource for graph reasoning research. The detailed comparison with previous related benchmarks is shown in Table 2.
3.2. Supervised Training of GraphPRM
PRMs evaluate whether a reasoning subsequence is on the correct path by assigning a correctness score. The PRM is trained via supervised fine-tuning on an LLM using correct/incorrect labels. PRM800K (Lightman et al., 2024) classifies each step as ’good’, ’neutral’, or ’bad’. However, for graph reasoning tasks, the difference between binary and three-class classification is minimal, so we treat PRM training as a binary classification task.
[…]
(3) $\min_{\theta}\mathcal{L}(\theta)=-\frac{1}{N}\sum_{t=1}^{T}m_{t}\Big{[}y_{t}\log\hat{y}_{t}+(1-y_{t})\log(1-\hat{y}_{t})\Big{]},$
where $\theta$ is the parameter of GraphPRM, $y_{i}$ is the $i$-th reasoning step label and $\hat{y}_{i}=P(y_{i}\mid p,s_{1},\dots,s_{i})$ is the predicted probability of $i$-th step. $m_{t}$ indicates whether each token is the step token.
3.3. Scaling Inference-Time Compute via PRMs
3.3.1. Inference-time Searching Strategies
Following prior work (Lightman et al., 2024; Snell et al., 2024), we use GraphPRM to evaluate individual steps in solutions from the policy model. We explore two paradigms: (1) Best-of-N selection and (2) Guided beam search, integrating PRM for guided decoding and scoring or voting across solutions.
Best-of-N Selection The Best-of-N search strategy involves sampling $N$ candidate solutions independently from the base LLM and selecting the most optimal solution based on a PRM’s evaluation. Specifically, for a given problem $p$ in the test set, $N$ candidate solutions $s^{*}$ are generated by the LLM, and each step $s_{i}$ in solution $s^{*}$ is scored by GraphPRM. With answer aggregation approach, we can assess the quality of each solution. The solution with the highest score is chosen as the final answer. By leveraging GraphPRM, this approach increases the likelihood of selecting a correct solution, thereby enhancing the success rate of LLMs in solving GCPs.
Guided Beam Search Guided beam search uses PRM for step-by-step optimization. Following BFS-V (Yao et al., 2023), the LLM generates $N$ candidates for the first step, which are then evaluated and scored by the PRM. The top $K(K\leq N)$ candidates with the highest scores are retained for the current step. For each of these $K$ outputs, the base LLM generates $M=N/K$ (where $M\in\mathbb{Z}$) subsequent candidates, reconstituting a total pool of $N$ outputs. This iterative process continues, with new candidates being scored, filtered, and sampled at each step. Crucially, the PRM’s scores serve as the primary mechanism for guiding and refining the search trajectory at each step.
3.3.2. Answer Aggregation
During inference, GraphPRM scores each step of a solution from the base LLM. To identify the final answer, a two-step aggregation function is required to aggregate per-step scores to determine the most likely correct solution. First, we combine the per-step scores to compute a final score for the entire single solution. Following (Lightman et al., 2024; Snell et al., 2024), we adopt two approaches to aggregate step-wise scores for a single solution $s$: 1) PRM-Last: We take the score of the last step $s_{e}$ as the full-solution score, denoted as $v_{s^{*}}=r_{s_{e}}$. 2) PRM-Min: Choose the minimum score among all steps to be the overall score, denoting as $v_{s^{*}}=\min\left(r_{s_{1}},r_{s_{2}},\dots,r_{s_{e}}\right)$, which is demonstrated as good as PRM-Last strategy (Snell et al., 2024). After that, we aggregate the final scores across all solutions. Through a Weighted Majority Vote, combining self-consistency and PRM, solutions are grouped by their final answers, and an aggregate score is computed for each group to select the best overall answer. Formally, the final answer based on $N$ candidate solutions is:
[…]
where $z_{i}$ represents the answer of $i$-th generated solution, and $v_{s^{*}}$ denotes the aggregated score of solution $s^{*}$, which can be implemented as either PRM-Last or PRM-Min. Notably, we denote the above combined two multi-answer weighting methods as PRM-Last-Vote and PRM-Min-Vote.
Verification Strategy In-domain Tasks Average Out-of-domain Tasks Average
Degree Clustering Jaccard Common Connectivity Diameter Page Rank MST Flow Predecessor Neighbor BFS Cycle
Qwen2.5-1.5B-Instruct
Self-Consistency 84.33 9.00 32.33 44.67 31.00 13.33 8.33 0.33 6.33 3.67 23.33 33.33 8.33 50.33 30.67
GraphPRM (Best-of-N) 86.67 9.00 37.00 45.67 32.67 12.33 10.00 0.33 5.67 3.33 24.27 39.67 9.67 50.33 33.22
GraphPRM (Beam Search) 92.67 13.00 53.00 45.00 67.33 15.00 10.67 1.00 3.67 2.33 30.37 52.33 7.00 51.00 36.78
Qwen2.5-7B-Instruct
Self-Consistency 90.00 25.33 58.00 49.00 43.33 48.00 12.33 0.33 13.33 30.67 37.03 69.67 48.00 52.33 56.67
GraphPRM (Best-of-N) 90.67 26.33 60.00 49.33 49.67 46.67 13.67 0.33 14.00 32.33 38.30 73.33 49.00 53.00 58.44
GraphPRM (Beam Search) 94.33 22.33 74.33 48.00 79.33 43.33 11.33 0.00 7.33 35.33 41.57 86.00 33.67 53.33 57.67
LLaMA3.1-8B-Instruct
Self-Consistency 84.00 10.00 44.00 46.33 23.00 17.33 15.33 0.00 6.67 12.67 25.93 61.00 26.00 49.67 45.56
GraphPRM (Best-of-N) 87.00 12.67 43.33 46.33 24.00 18.00 16.33 0.33 7.00 12.33 26.73 65.00 26.67 50.00 47.22
GraphPRM (Beam Search) 94.33 14.33 62.33 47.33 37.00 26.33 9.33 1.00 7.33 14.00 31.33 78.00 18.33 38.33 44.89
Gemma2-9B-Instruct
Self-Consistency 95.67 10.00 41.00 33.33 55.00 31.67 13.00 0.00 7.67 30.33 31.77 93.67 31.67 54.33 59.89
GraphPRM (Best-of-N) 97.00 10.67 41.67 34.00 59.33 33.67 12.67 0.00 7.67 30.67 32.73 94.33 35.67 59.67 63.22
GraphPRM (Beam Search) 97.67 12.67 43.00 39.00 77.67 25.33 9.67 0.00 8.67 28.00 34.17 90.00 35.33 56.67 60.67
Table 3. Performances (%) of different LLMs on GraphSilo in both in-domain and out-of-domain tasks with different verification strategies. The reward model is trained with Qwen2.5-Math-7B on GraphSilo. For both Best-of-N and Beam Search strategies, the generation budget is set to 8 and the answer aggregation strategy is PRM-Last. The best results are in Purple.
3.4. RL Training with Process Supervision
Upon achieving GraphPRM, we employ reinforcement learning to further train LLMs. Considering the instability of PPO (Schulman et al., 2017) training, we choose Direct Preference Optimization (DPO) (Rafailov et al., 2023) instead. Under this circumstance, GraphPRM guides the selection of high-quality preference pairs, ensuring correctness in preferred solutions and falsehood in dispreferred ones. Thus, DPO refines the original policy model by aligning its parameter $\pi(\theta)$ more closely with preferred outcomes using a training corpus similar to the distribution from the policy model. To be specific, DPO employs input pairs labeled $(\mathcal{S}_{w},\mathcal{S}_{l})$, where $\mathcal{S}_{w}$ and $\mathcal{S}_{l}$ are the set of preferred and less preferred reasoning processes. To obtain these training data, we first employ Supervised Fine-Tuning (SFT) to train a base LLM with language modeling loss on all tasks from GraphSilo. The model is trained to process graph $\mathcal{G}$ and problem $\mathcal{P}$ as input and generate reasoning process $S$ as output, obtaining model $\mathcal{M}(\theta)_{sft}$ as follows:
(5) $\begin{array}[]{l}\mathcal{L}_{\mathrm{LM}}=-\sum_{i=1}^{N}\sum_{j=1}^{M}\log\mathrm{P}\left(\mathcal{S}_{i,j}\mid\mathcal{G}_{i},P_{i};\theta\right),\end{array}$
where $S$ is the set of reasoning processes towards graph problems, $N$ denotes the number of problems and $M$ represents the number of reasoning processes for each graph problem. Then we generate a new set of 20,000 graph problems, and perform two strategies on $\mathcal{M}(\theta)_{sft}$ to obtain the preference dataset: 1) By performing Beam Search with 8 width on $\mathcal{M}(\theta)_{sft}$ with our trained GraphPRM, we take the best output of each problem as its preferred reasoning process $\mathcal{S}_{w}$. 2) We apply Best-of-N on $\mathcal{M}(\theta)_{sft}$ with GraphPRM, generating 8 samples per problem. The sample with the wrong answer and lowest step score is chosen as the dispreferred reasoning process $\mathcal{S}_{l}$. Through this we can obtain fine-grained preference dataset $\mathcal{D}$ with process supervision, resulting in 4,716 $(\mathcal{S}_{w},\mathcal{S}_{l})$ pairs. The formal training objective for DPO is defined as follows:
(6) $\begin{array}[]{l}\mathcal{L}_{DPO}\left(\mathcal{M}(\theta);\mathcal{M}(\theta)_{\mathrm{sft}}\right)=-\mathbb{E}_{\left(x,\mathcal{S}_{w},\mathcal{S}_{l}\right)\sim D}\\ {\left[\log\sigma\left(\beta\log\frac{\mathcal{M}(\theta)\left(\mathcal{S}_{w}\mid x\right)}{\mathcal{M}(\theta)_{\mathrm{sft}}\left(\mathcal{S}_{w}\mid x\right)}-\beta\log\frac{\mathcal{M}(\theta)\left(\mathcal{S}_{l}\mid x\right)}{\mathcal{M}(\theta)_{\mathrm{sft}}\left(\mathcal{S}_{l}\mid x\right)}\right)\right],}\end{array}$
where $x$ is the concatenation of problem and reasoning processes and $\beta$ is the preference hyper-parameter.
4. Experiments
In this section, we evaluate the proposed GraphPRM by addressing the following research questions:
  • RQ1: How does GraphPRM scales inference-time performance over LLMs on GraphSilo, particular when transferring to unseen and out-of-dataset graph tasks?
  • RQ2: How transferable is GraphPRM to mathematical reasoning problems?
  • RQ3: How does GraphPRM benefit LLMs reasoning abilities via Preference Alignment with process supervision?
  • RQ4: How does graph size affect the performance of GraphPRM on LLMs during inference-time?
  • RQ5: How does process label quality affect GraphPRM?
  • RQ6: What is the influence of parameter sizes on the interplay between policy model and GraphPRM?
4.1. Experimental Settings
Datasets We conduct experiments on both graph reasoning and math reasoning tasks. For graph reasoning, we use our GraphSilo alongside the widely used GraphWiz (Chen et al., 2024a) as the test set. Tasks in GraphSilo are divided into In-domain and Out-of-domain groups based on their presence in GraphPRM’s training data. We designate BFS, Cycle, and Neighbors as Out-of-domain tasks because: 1) These three tasks are relatively challenging, requiring LLMs to possess a certain level of reasoning abilities to complete; 2) They cover both node-level and graph-level aspects, and outputs involve diverse formats. For math reasoning, we evaluate widely-used datasets, including GSM8K (Cobbe et al., 2021a), GSM8K-Hard (Gao et al., 2023), MATH500 (Hendrycks et al., 2021), and SVAMP (Patel et al., 2021). We evaluate the LLM’s output using Exact-Match accuracy (Snell et al., 2024). For float outputs, a relative error of less than 1e-4 is required. Evaluations use a zero-shot prompt setting, with results reported as mean accuracy over three sampled groups. In the reinforcement learning setting, the entire test set is used to evaluate model performance.
Setting on Inference-time Scaling Experiments are based on a series of large language models including Qwen2.5-1.5B/7B, Qwen2.5-Math-1.5B/7B (Team, 2024b), LLaMA3.1-8B (AI@Meta, 2024) and Gemma2-9B(Team, 2024a). We train our GraphPRM on Qwen2.5-Math-7B-Instruct with 118k step-level data from GraphSilo. GraphPRM is trained for three epochs with a learning rate of 1e-4. During inference-time computation, we set the decode number $N$ to 8 and temperature parameter to 0.7 for both searching strategies in most experiments. Among different aggregation strategies, we select PRM-Last-Vote as a representative. The LLM inference server is implemented using FastChat (Zheng et al., 2023; Wang et al., 2024a).
Setting on Reinforcement Learning In the reinforcement learning experiments, we use the Qwen2.5-7B-Instruct model as the base model for SFT and DPO training, with the trained GraphPRM serving as the PRM to select and construct the preference dataset for DPO training. For SFT training, we construct a new set of graph tasks in GraphSilo containing 20,000 problem-process pairs. In DPO training, the learning rate is 5e-6, and the preference hyperparameter $\beta$ is set to 0.1.
Figure 4. Best-of-N Performance of GraphPRM on four math datasets with ranging generation budget (number of solution candidates). The policy model is Qwen2.5-Math-1.5B-Instruct and the answer aggregation strategy is PRM-Last.
4.2. Performance of GraphPRM (RQ1)
To evaluate GraphPRM, we conduct inference-time scaling experiments with Best-of-N and Beam Search strategies on GraphSilo, and performance is reported in Table 3. Notably, compared to Self Consistency, LLMs with GraphPRM achieve remarkable improvements during inference-time computation with the same generation budget, especially under the Beam Search strategy via GraphPRM. This consistency is maintained across a range of tasks at different levels. Besides, GraphPRM achieves a general better performance transferring to those out-of-domain tasks (Neighbor, BFS, Cycle) that are not included in the training dataset of GraphPRM.
Dataset Transfer In addition to our GraphSilo, we also evaluate GraphPRM on the widely used GraphWiz dataset. We select five most representative tasks (Connect, Topology, Shortest, Hamilton, and Subgraph) across three different difficulty levels and directly test GraphPRM without using example prompts or additional training. As shown in Table 4, GraphPRM demonstrates significant improvements in LLMs through inference-time computation, highlighting its effectiveness across datasets with diverse distributions.
Verification Strategy Easy Medium Hard Average
Connect Topology shortest Hamilton Subgraph
Qwen2.5-7B-Instruct
Self-Consistency 29.75 15.00 18.00 0.25 9.50 14.50
GraphPRM (Best-of-N) 32.50 13.25 17.00 10.00 25.25 15.80
GraphPRM (Beam Search) 34.50 8.25 14.00 29.75 14.00 20.10
LLaMA3.1-8B-Instruct
Self-Consistency 57.00 6.50 8.25 42.25 41.00 31.00
GraphPRM (Best-of-N) 60.50 7.25 8.50 44.50 43.50 32.85
GraphPRM (Beam Search) 82.75 12.25 10.50 51.25 57.25 42.80
Table 4. Performances (%) of GraphPRM on GraphWiz dataset in five representative tasks. The generation budget during verification is set to 8. The best results are in Purple.
Generation Budget We further evaluate GraphPRM on LLMs with varying numbers of generation samples to analyze the impact of the generation budget (representing computational effort or token usage per question) on inference-time scaling. As shown in Figure 5, both the Best-of-N and Beam-Search methods significantly outperform Self Consistency, particularly as the generation budget increases. This confirms that GraphPRM enhances LLM performance during inference-time computation, with improvements closely tied to the generation budget.
Figure 5. Performance of GraphPRM with different evaluation paradigms. The generation budget ranges from 1 to 128.
4.3. Transferability on Math Problems (RQ2)
To investigate the zero-shot transferability of GraphPRM on mathematical reasoning tasks, we conduct experiments on four classic math datasets. Figure 4 illustrates the performance under the Best-of-N strategy during inference-time computation with varying generation budgets. It is clear that GraphPRM significantly enhances the performance of the policy model on mathematical reasoning tasks and consistently outperforms the Self-Consistency method across all generation budgets. Besides, the performance gap between GraphPRM and Self Consistency is more pronounced in datasets with higher complexity (e.g., GSM8K Hard and MATH) compared to simpler datasets (e.g., SVAMP).
Regarding the phenomenon that GraphPRM achieves significant zero-shot transferability on mathematical reasoning tasks, we attribute it to the following reasons: 1) Graph algorithms are fundamentally grounded in mathematical concepts, such as graph theory and combinatorics, requiring models to understand complex relational modeling and optimization logic, which align closely with the demands of mathematical reasoning. 2) The abstraction inherent in graph algorithms—dealing with nodes, edges, and their relationships—enhances the model’s capacity to process variables and sets in mathematical reasoning. Additionally, the recursive and divide-and-conquer strategies learned in graph algorithms (e.g., depth-first search or dynamic programming) can be directly applied to mathematical reasoning tasks, such as solving recursive relations or constructing complex proofs. 3) Graph algorithm tasks often involve dynamic inputs (e.g., variable graph structures, changing edge weights), requiring models to flexibly adjust their reasoning processes. This dynamic adaptability is equally important when dealing with open-ended and variable mathematical problems.
4.4. Performance of Post-trained LLMs with Process Supervision (RQ3)
We report the performance of post-trained LLMs with process supervision via GraphPRM in Table 5. ”DPO-Outcome” refers to a model post-trained on a preference dataset where preferred and dispreferred reasoning processes for each problem are randomly selected based on their outcome answers (e.g., a reasoning process with a correct answer is randomly selected from all generated candidates as the preferred one). We perform DPO training on 1) vanilla Qwen2.5-7B model and 2) Qwen2.5-7B model supervised fine-tuned on all 9 training tasks from GraphSilo, and results indicate that DPO with process supervision via GraphPRM significantly enhances the performance of two policy models, achieving improvements of 12.4% and 10.91%, respectively, and outperforming the ”DPO-Outcome” ablation. Furthermore, integrating reinforcement learning with inference-time scaling using GraphPRM proves complementary, delivering an additional 1% performance boost.
Models GraphSilo
Qwen2.5-7B-Instruct 34.60
+ DPO - Outcome 35.47
+ DPO - Process Supervision (Ours) 43.00
+ DPO - Process Supervision + GraphPRM (Best-of-8) 44.20
Qwen2.5-7B-Instruct: SFT 53.57
+ DPO - Outcome 55.77
+ DPO - Process Supervision (Ours) 58.47
+ DPO - Process Supervision + GraphPRM (Best-of-8) 59.03
Table 5. Performances of reinforcement learning and verification combination. The best results are in Purple.
4.5. Results with Increasing Graph Size (RQ4)
To further investigate how graph sizes of test graph problems affect the performance of GraphPRM on backbone LLMs, we generate four different test sets in various sizes: Tiny, Small, Medium and Large, each with 300 problems. Results on these four test sets are shown in Figure 6. It can be shown that for all inference-time searching strategies, accuracy declines with a decrease percentage around 53% as graph size increases. Self Consistency starts with relatively high accuracy for Tiny graphs but shows a sharp decline for Medium and Large graphs. On the other hand, GraphPRM with Best-of-N and Beam Search methods exhibit greater robustness, especially for Small and Medium graphs, maintaining relatively higher and more consistent accuracy compared to Self Consistency. Comparing performance among three methods on each graph size, GraphPRM consistently performs beyond Self Consistency.
Beyond this, it is evident that with the increase of graph size, GraphPRM have more improvement on policy LLMs under Beam Search strategy, while have less improvement on policy LLMs under Best-of-N strategy. This decline can likely be attributed to two main factors: 1) As graph size increases, the complexity of the problem grows, the probability of the fixed candidate pool containing optimal or near-optimal solutions decreases, making it harder for Best-of-N to identify optimal solutions. 2) For larger graphs, the search space expands significantly, and the complexity of finding the best solution grows. Beam Search, supported by GraphPRM, can better adapt to this complexity by incorporating voting mechanisms to systematically narrow down the search space. This observation underscores the effectiveness and robustness of GraphPRM over different sizes of graphs.
Figure 6. Performance of GraphPRM on GraphSilo with various graph sizes. The policy model is Qwen2.5-7B-Instruct and the verification is based on 8 candidates.
4.6. Effect of Different Process Labeling Strategies (RQ5)
To better understand the effectiveness of our proposed process annotation approaches discussed in Section 3.1.3, we perform an ablation study on the quality of process annotations in GraphSilo. As shown in Table 6, GraphPRM benefits significantly from both types of high-quality process annotations. When comparing the two annotation approaches, task-oriented labels yield better results under the Best-of-N searching strategy, achieving a 2.5% improvement, whereas Monte-Carlo labels perform better under the Beam Search strategy, with a 7.2% improvement. We identify the main rationales behind this phenomenon are: 1) Task-oriented labels are carefully designed to match the graph problem’s structure and provide clear, step-by-step logical guidance. It aligns well with the static, fixed-pool nature of Best-of-N; 2) Monte-Carlo labels excel in the Beam Search strategy because their diversity and variability allow the iterative refinement process to explore a wider solution space, maximizing the potential for improvement; 3) Combining both comprehensive reasoning processes with fine-grained step-level labels from these two approaches further enhance the effectiveness of GraphPRM.
Methods Best-of-N Beam Search
Qwen2.5-7B-Instruct (Self Consistency) 37.03 37.03
+ GraphPRM 38.30 41.57
+ GraphPRM - Task-oriented Labels 37.96 36.03
+ GraphPRM - Monte-carlo Labels 37.36 39.70
Table 6. Performances (%) of GraphPRM trained on different set of step-wise process labels. The verification is based on 8 candidates. The best results are in Purple.
4.7. Influence of Pre-trained Model Sizes (RQ6)
Figure 7. Best-of-N Performance of GraphPRM on GraphSilo with policy models and reward models in different parameter sizes. The answer aggregation strategy is PRM-Last.
To conduct an exhaustive analysis of to what extent does the combination of model parameters affect the final performance on graph reasoning tasks, we performed a range of experiments using model sizes in 1.5B and 7B. Figures 7(a) and 7(d) display the results from the 1.5B and 7B policy models paired with equal-sized GraphPRMs, respectively. It becomes evident that GraphPRM exhibits superiority over Self Consistency across all sizes of base models.
Figure 7(b) and 7(c) presents the performance of 1.5B and 7B policy models interfaced with different-sized GraphPRMs. The findings illustrate that utilizing a larger GraphPRM to validate the output of a smaller policy model significantly enhances performance. Conversely, when a smaller GraphPRM is employed to validate the output of a larger policy model, the verification process adversely impacts the model’s performance compared to Self Consistency. These results substantiate that we should utilize a more potent GraphPRM for validating or supervising the policy model.
4.8. Case Study
In this subsection, we present two detailed cases on both graph reasoning problem and mathematical reasoning problem, in which we compare the correct and wrong reasoning processes generated from policy LLM model, alongside with scores assigned by GraphPRM.
Case on Graph Reasoning Figure 8 shows a case study of a Clustering Coefficient problem in GraphSilo dataset. It can be observed that when presented with a graph reasoning problem, our GraphPRM accurately selected the correct solution from a pool of candidate solutions. Specifically, in wrong solution case, LLM starts fall in wrong in step 2 when calculating the neighbors of node 5 (regarding 6 as its neighbor), thus leading to subsequent mistakes in reasoning steps. Our GraphPRM displays superior discernment by precisely identifying wrong steps and assigns low scores from step 2 to step 7. In correct solution case, GraphPRM precisely rank high scores to correct steps. This case study further demonstrate the effectiveness of GraphPRM.
Case on Mathematical Reasoning Figure 9 presents a case study of a mathematical reasoning problem from the GSM8K dataset. The results in this figure highlight the effectiveness of GraphPRM in verifying the correctness of the mathematical reasoning process, ultimately yielding a correct final answer through reliable step-by-step reasoning. Notably, we observe that the LLM begins to falter in step 3, where it calculates the profit for the jewelry purchase. At this step, it incorrectly interprets ”the total value of the electronic gadget” as the ”profit”, despite performing the calculation in this step correctly. GraphPRM successfully identifies this misunderstanding of the calculation target and assigns a low score to this step, along with subsequent steps that depend on the erroneous reasoning from step 3. Conversely, in the correct solution, GraphPRM assigns high scores to the accurate steps. This case study further demonstrates the adaptability and transferability of GraphPRM for mathematical reasoning tasks.
Figure 8. A case of Clustering Coefficient task in GraphSilo. Correct and wrong solutions are identified with GraphPRM via Best-of-N strategy.
Figure 9. A case of mathematical reasoning result in GSM8K dataset. Correct and wrong solutions are identified with GraphPRM via Best-of-N strategy.
5. Conclusion
In this paper, we introduce a novel approach to enhance reasoning capabilities in LLMs by leveraging PRMs for graph reasoning tasks. We developed GraphSilo, the largest dataset for graph reasoning with fine-grained, step-wise labels. Building upon this dataset, we proposed GraphPRM, the first PRM tailored for reasoning on GCPs, and demonstrated its effectiveness through extensive evaluations in settings such as inference-time scaling and reinforcement learning via Direct Preference Optimization (DPO). Our experiments show that GraphPRM delivers significant improvements across 13 graph reasoning tasks and demonstrates strong generalization to unseen graph reasoning datasets. Notably, we observed a compelling transferability of GraphPRM to other reasoning domains, including mathematical problem-solving tasks like GSM8K and MATH500. These results underscore the potential of PRMs to bridge reasoning domains and highlight their promise in advancing step-wise reasoning, laying the groundwork for applying graph-based reasoning to broader challenges and advancing LLM capabilities.
Appendix A More Dataset Details
A.1. Graph Structure Details
In Section 3.1.2, we elaborate on the diverse graph structure distributions employed in the construction of GraphSilo. Specifically, we incorporate three fundamental network topologies:
  • Random networks: These networks are constructed by establishing edges between $N$ nodes with probability $p$, resulting in a stochastic connectivity pattern.
  • Scale-free networks: Characterized by a power-law degree distribution, these networks are prevalent in various real-world systems, including social interactions, biological processes, and economic trade relationships.
  • Small-world networks: These networks exhibit a distinctive property where, despite relatively sparse direct connections, any two nodes can typically be connected through a small number of intermediate nodes, facilitating efficient information flow across the network.
A.2. Dataset Statistics
Table 7 presents the dataset statistics for GraphSilo, providing detailed information on the PRM Training Set, PRM Test Set, SFT Training Set, and DPO Training Set for better understanding.
Property Number
PRM Training Set
# of graph tasks 13
# of problems 20,063
# of problem-solution pairs (total) 118,189
# of problem-solution pairs (Task-oriented Trajectory) 75,852
# of problem-solution pairs (Monte-carlo Method) 42,337
# of positive step labels 220,894
# of negative step labels 173,271
# of maximum steps in a solution 20
PRM Test Set
# of problems (total) 3,900
# of problems per task 300
# of graph size types 4
SFT Training Set
# of problem-solution pairs (total) 7,964
DPO Training Set
# of positive-negative preference pairs (origin Qwen) 6,137
# of positive-negative preference pairs (SFT model) 5,031
Table 7. Data Statistics of GraphSilo
A.3. Quality Control for GraphSilo
Ideally step-wise process labels should be clear enough to support the validity of corresponding solution sub-steps. After obtaining graph reasoning processes with step-level labels via approaches proposed in Section 3.1.3, we conduct quality control pipeline to select high-quality data for further analysis. Specifically, we include the following three steps:
Completeness Filtering All reasoning processes in GraphSilo should be meaningful and related to corresponding graph problem. For data generated from ”Task-oriented trajectories” approach, their completeness and correctness are inherently assured due to the advantages of graph algorithms. For data generated from ”Automated Monte-Carlo Method” approach, we filter out those: 1) reasoning processes without a final answer; 2) reasoning processes containing repeated meaningless sentences.
Label Filtering Each step in a solution is assigned a binary label, ‘+’ for positive and ‘-’ for negative. For problem-solution pairs generated using the ”Task-oriented Trajectories” method, we retain only those pairs where negative steps occur exclusively after positive steps. For pairs generated via the ”Automated Monte-Carlo Method,” all steps following an incorrect step (label ‘-’) are removed, as the validity of correct step becomes irrelevant after an error. This removal is designed to avoid potential confusion during model training (Zheng et al., 2024).
Human Validation We recruited five volunteers to evaluate the entailment between step-wise label and corresponding reasoning sub-step. Each volunteer assessed a sample of 500 randomly selected examples to ensure the quality and reliability of our dataset. They were tasked with ”Determining whether the step-wise label supports the corresponding reasoning step regrading the graph problem.” The human evaluation results are listed in Table 8. The high agreement observed further supports our dataset’s quality and validity of step-wise process labels.
Anno-1 Anno-2 Anno-3 Anno-4 Anno-5
Agreement Rate 95.0 94.2 95.4 96.0 94.4
Table 8. Human evaluation results on GraphSilo
A.4. Detailed Task Definition
In this section, we present detailed definitions for each task included in the GraphSilo dataset. For tasks in Boolean output format, we curate a balanced dataset by filtering the graphs to ensure an equal distribution of positive and negative labels.
  • Task 1: Degree. Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, which may be either directed or undirected. For a specified node $v\in\mathcal{V}$, compute its degree: in an undirected graph, the degree is the number of incident edges; in a directed graph, determine both the in-degree (number of incoming edges) and the out-degree (number of outgoing edges).
  • Task 2: Clustering Coefficient. Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, which may be directed or undirected. For a specified node $u\in\mathcal{V}$, the task is to compute its clustering coefficient as follows:
    • For a directed graph, let $T$ be the number of edges among the out-neighbors of $u$ and $D$ be the out-degree of $u$. The clustering coefficient is then given by $C=\frac{T}{D(D-1)}$.
    • For an undirected graph, using the same definitions for $T$ and $D$, the clustering coefficient is defined as $C=\frac{2T}{D(D-1)}$.
    The objective is to accurately calculate and report the clustering coefficient value for node $u$ based on the type of graph provided.
  • Task 3: Neighbor. Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ that may be either directed or undirected. For a given node $u\in\mathcal{V}$, identify its neighbors—that is, all nodes directly connected to $u$. In an undirected graph, these are all nodes $v$ with ${u,v}\in\mathcal{E}$; in a directed graph, they include both nodes $v$ with $(u,v)\in\mathcal{E}$ (out-neighbors) and nodes $v$ with $(v,u)\in\mathcal{E}$ (in-neighbors).
  • Task 4: Page Rank. Compute the PageRank values for each node in graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ (directed or undirected) by initializing each node’s PageRank to $1/|\mathcal{V}|$, then performing 3 iterations of the algorithm with a damping factor of 0.85. Finally, identify the node with the largest PageRank value.
  • Task 5: Predecessor. Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ that may be either directed or undirected. For a given node $u$, return its predecessor nodes as follows: in a directed graph, a predecessor of $u$ is any node $m$ such that there is an edge from $m$ to $u$; in an undirected graph, treat all neighbors of $u$ as its predecessors.
  • Task 6: Jaccard. Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ that may be either directed or undirected. For two given nodes $u$ and $v$, compute their Jaccard similarity, defined as the size of the intersection of their neighbor sets divided by the size of the union of their neighbor sets. In an undirected graph, the neighbor set of a node consists of all nodes directly connected to it, while in a directed graph, it includes both in- and out-neighbors. Return the computed Jaccard similarity value for the pair $(u,v)$.
  • Task 7: Common Neighbor. Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ that may be either directed or undirected. For two given nodes $u$ and $v$, identify the common neighbors—that is, the nodes that are connected to both $u$ and $v$. In a directed graph, treat the union of in-neighbors and out-neighbors as the neighbor set for each node.
  • Task 8: Connectivity. Given graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ that may be either directed or undirected. the task is to assess if two randomly chosen nodes $u$ and $v$ are connected through a sequence of edges.
  • Task 9: Maximum Flow. Consider a directed, weighted graph $\mathcal{G}=\{\mathcal{V},\mathcal{E},c\}$, where $c:\mathcal{E}\to\mathbb{R}^{+}$ is a function assigning a positive capacity to each edge, representing the maximum flow that the edge can support. Given a source node $v_{s}$ and a sink node $v_{t}$ in $\mathcal{G}$, the task is to devise a plan to maximize the flow from the source $s$ to the sink $t$.
  • Task 10: Breadth First Search. Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ that may be either directed or undirected. For a given starting node $u$, perform a BFS traversal and return the sequence of nodes in the order they are visited. In a directed graph, use out-neighbors as adjacent nodes; in an undirected graph, use all directly connected neighbors.
  • Task 11: Cycle. In a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, the task is to detect the existence of a cycle. A cycle can be defined as a sequence of vertices $v_{1},v_{2},\dots,v_{k}$ with $k\geq 3$, that forms a closed loop, meaning $v_{1}=v_{k}$. Additionally, for all $1\leq i<k$, each vertex $v_{i}$ must be distinct from the others, and there must be an edge connecting $v_{i}$ to $v_{i+1}$.
  • Task 12: Diameter. Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, which may be either directed or undirected. The task is to compute the diameter of $\mathcal{G}$, defined as the maximum shortest path distance between any two nodes. Return the computed diameter.
  • Task 13: Minimum Spanning Tree. Consider an undirected, weighted graph $\mathcal{G}=(\mathcal{V},\mathcal{E},w)$, where $w$ assigns a weight to each edge. The task is to compute the Minimum Spanning Tree (MST) of $\mathcal{G}$—a subset of edges that connects all nodes with the minimum total edge weight. Return the MST or its total weight.
Appendix B Additional Experimental Results
In this section, to further analysis our proposed GraphPRM, we evaluate GraphPRM by addressing the following question: ”How does GraphPRM perform on other GCP-specific backbones?”
Performance of GraphPRM on GraphWiz model In addition to our proposed GraphSilo, we conduct experiments on GraphWiz dataset with Qwen2.5-7B-Instruct, Qwen2.5-Math-7B-Instruct and LLaMA-3.1-8B-Instruct in Section 4.2. Besides above, we further evaluate the transferbility of GraphPRM on other GCP-specific model like GraphWiz. As shown in Figure 10, GraphPRM consistently improve LLMs’ performance on GraphWiz dataset with more than 5% improvement.
Figure 10. Best-of-N Performance of GraphPRM on GraphWiz with different backbones. We adopt GraphWiz-DPO trained on LLaMA2-7B. Verification is based on 8 candidates.
Appendix C More Cases
C.1. Task Prompts Used in Experiments
In this section, we provide a detailed list of all prompts for all experiments, offering a clear reference for understanding our experimental approach. Prompts for evaluating graph reasoning tasks in GraphSilo on both inference-time computation and reinforcement learning settings are listed in Table 9.
Task Prompt
[…]
Table 9. Prompts for evaluating graph reasoning tasks in GraphSilo on both inference-time computation and reinforcement learning settings. For readability, we omit the input graph edge list in the format of natural language.
C.2. Cases with GraphPRM Scores
In this section, we provide a detailed list of reasoning cases scored by GraphPRM, offering a clear reference for understanding the effectiveness of GraphPRM:
  • Cases of reasoning results on Connectivity, Predecessor and Diameter tasks in GraphSilo are listed in Figure 11, Figure 12 and Figure 13, respectively.
  • Cases of mathematical reasoning results on MATH dataset is listed in Figure 14.
Figure 11. A case of Connectivity task in GraphSilo. Correct and wrong solutions are identified with GraphPRM via Best-of-N strategy.
Figure 12. A case of Predecessor task in GraphSilo. The best solution is identified with GraphPRM via Beam Search strategy.
Figure 13. A case of Diameter task in GraphSilo. The best solution is identified with GraphPRM via Beam Search strategy.
Figure 14. A case of mathematical reasoning result in MATH dataset. Correct and wrong solutions are identified with GraphPRM via Best-of-N strategy.

### v2

@@
 Our findings highlight the potential of PRMs in advancing reasoning across diverse domains, paving the way for more versatile and effective LLMs.
-Models and data are available at https://huggingface.co/GraphPRM.
+Codes and data are available at https://github.com/GKNL/GraphPRM.
 Large Language Models, Graph Reasoning, Process Reward Model
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Natural language processing
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737109isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Natural language processing
+KDD Availability Link:
+The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.15553493.
 1. Introduction
@@
 This enables us to explore automated solution paths for GCPs while utilizing reward models to refine the reasoning process.
+Level Task Definition Time Complexity Answer Type
+Node # Degree Return the degree of node $u$ in a given graph $\mathcal{G}$. $O(|\mathcal{V}|)$ Integer
+# Clustering Coefficient Calculate the ratio of actual edges to possible edges among node $u$’s neighbors in a graph $\mathcal{G}$. $O(|\mathcal{V}|^{2})$ Float
+# Neighbor Find the neighbors of node $u$ in a given graph $\mathcal{G}$. $O(|\mathcal{V}|)$ Node List
+# Page Rank Identify the node with the highest PageRank value in a given graph $\mathcal{G}$. $O(|\mathcal{V}|+|\mathcal{E}|)$ Integer
+# Predecessor Return the list of predecessor nodes of $u$ in a given graph $\mathcal{G}$ $O(|\mathcal{V}|)$ Node List
+# Jaccard Calculate the Jaccard Coefficient value $J$ between node $u$ and node $v$, where $J=\frac{|\mathcal{N}(u)\cap\mathcal{N}(v)|}{|\mathcal{N}(u)\cup\mathcal{N}(v)|}$. $O(|\mathcal{V}|)$ Float
+Node # Common Neighbor Find the number of common neighbors of node $u$ and node $v$ in a given graph $\mathcal{G}$. $O(|\mathcal{V}|)$ Integer
+Pair # Connectivity Determine if there is a path from node $u$ to node $v$ in a given in a given graph $\mathcal{G}$. $O(|\mathcal{V}|+|\mathcal{E}|)$ Boolean
+# Maximum Flow Calculate the maximum amount of flow from node $u$ to node $v$ in a given graph $\mathcal{G}$. $O(|\mathcal{V}|^{2}\sqrt{|\mathcal{E}|})$ Integer
+Graph # Breadth First Search Return the sequence of nodes in the order of a BFS traversal starting from node $u$. $O(|\mathcal{V}|+|\mathcal{E}|)$ Node List
+# Cycle Determine if there exists an cycle in a given graph $\mathcal{G}$. $O(|\mathcal{V}|+|\mathcal{E}|)$ Boolean
+# Diameter Calculate the maximum distance over all pairs of nodes in a given graph $\mathcal{G}$. $O(|\mathcal{V}|(|\mathcal{V}|+|\mathcal{E}|)\text{log}|\mathcal{V}|)$ Integer
+# Minimum Spanning Tree Calculate the total weight of the minimum spanning tree in a given graph $\mathcal{G}$. $O(|\mathcal{E}|\text{log}|\mathcal{E}|)$ Integer
+Table 1.
+Descriptions of graph tasks in GraphSilo with problem definition, time complexity of representative algorithms and answer type.
+$|\mathcal{V}|$ and $|\mathcal{E}|$ indicate the number of nodes and edges in the graph $\mathcal{G}$.
 In light of these insights, this study explores the application of PRMs within GCP tasks to strengthen LLM.
@@
 Ultimately, our findings show that integrating graph-based reward models like GraphPRM makes LLMs more versatile and effective across various reasoning tasks, bridging gaps between graph computation and general reasoning applications.
-Figure 2.
-Overall illustration of process annotation pipeline of GraphSilo (a, b) and the strategies of utilizing GraphPRM to strengthen LLMs on graph reasoning, including scaling inference-time computation (c) and reinforcement learning (d).
 2. Preliminaries and Related Works
@@
 Reasoning - the capacity to derive new insights, tackle challenges, and reach conclusions from given information - stands as a core capability of LLMs.
-Researchers have proposed various methods to improve or elicit the reasoning ability of LLMs (e.g., mathematical reasoning), which can be broadly divided into three groups: pre-training (OpenAI, 2023; AI@Meta, 2024; Team, 2024b, a), fine-tuning (Yu et al., 2024; Luo et al., 2023; Yue et al., 2024; Chen et al., 2024c) and prompting (Imani et al., 2023; Wang et al., 2023c; Chen et al., 2024b).
+Researchers have proposed various methods to improve or elicit the reasoning ability of LLMs (e.g., mathematical reasoning), which can be broadly divided into three groups: pre-training (OpenAI, 2023; AI@Meta, 2024; Team, 2024b, a), fine-tuning (Yu et al., 2024; Luo et al., 2023; Yue et al., 2024; Chen et al., 2024c; Wu et al., 2025) and prompting (Imani et al., 2023; Wang et al., 2023c; Chen et al., 2024b).
 Except for directly improving and eliciting the reasoning potential of LLMs, researchers have been dedicated to boosting LLM reasoning results via an extra verifier for selecting the best answer from multiple decoded candidates.
@@
 The performance of the ORM is critically dependent on the construction of a high-quality training dataset with outcome labels.
-Level Task Definition Time Complexity Answer Type
-Node # Degree Return the degree of node $u$ in a given graph $\mathcal{G}$. $O(|\mathcal{V}|)$ Integer
-# Clustering Coefficient Calculate the ratio of actual edges to possible edges among node $u$’s neighbors in a graph $\mathcal{G}$. $O(|\mathcal{V}|^{2})$ Float
-# Neighbor Find the neighbors of node $u$ in a given graph $\mathcal{G}$. $O(|\mathcal{V}|)$ Node List
-# Page Rank Identify the node with the highest PageRank value in a given graph $\mathcal{G}$. $O(|\mathcal{V}|+|\mathcal{E}|)$ Integer
-# Predecessor Return the list of predecessor nodes of $u$ in a given graph $\mathcal{G}$ $O(|\mathcal{V}|)$ Node List
-# Jaccard Calculate the Jaccard Coefficient value $J$ between node $u$ and node $v$, where $J=\frac{|\mathcal{N}(u)\cap\mathcal{N}(v)|}{|\mathcal{N}(u)\cup\mathcal{N}(v)|}$. $O(|\mathcal{V}|)$ Float
-Node # Common Neighbor Find the number of common neighbors of node $u$ and node $v$ in a given graph $\mathcal{G}$. $O(|\mathcal{V}|)$ Integer
-Pair # Connectivity Determine if there is a path from node $u$ to node $v$ in a given in a given graph $\mathcal{G}$. $O(|\mathcal{V}|+|\mathcal{E}|)$ Boolean
-# Maximum Flow Calculate the maximum amount of flow from node $u$ to node $v$ in a given graph $\mathcal{G}$. $O(|\mathcal{V}|^{2}\sqrt{|\mathcal{E}|})$ Integer
-Graph # Breadth First Search Return the sequence of nodes in the order of a BFS traversal starting from node $u$. $O(|\mathcal{V}|+|\mathcal{E}|)$ Node List
-# Cycle Determine if there exists an cycle in a given graph $\mathcal{G}$. $O(|\mathcal{V}|+|\mathcal{E}|)$ Boolean
-# Diameter Calculate the maximum distance over all pairs of nodes in a given graph $\mathcal{G}$. $O(|\mathcal{V}|(|\mathcal{V}|+|\mathcal{E}|)\text{log}|\mathcal{V}|)$ Integer
-# Minimum Spanning Tree Calculate the total weight of the minimum spanning tree in a given graph $\mathcal{G}$. $O(|\mathcal{E}|\text{log}|\mathcal{E}|)$ Integer
-Table 1.
-Descriptions of graph tasks in GraphSilo with problem definition, time complexity of representative algorithms and answer type.
-$|\mathcal{V}|$ and $|\mathcal{E}|$ indicate the number of nodes and edges in the graph $\mathcal{G}$.
 Process Reward Model Different from ORM, PRM incorporates the concept of supervising intermediate reasoning steps in addition to the final outcome, enabling more granular feedback throughout the reasoning process.
@@
 To obtain high-quality process labels of PRM training dataset, existing methods include 1) human annotation (Lightman et al., 2024) and 2) automate annotation via Monte Carlo estimation (Wang et al., 2024b; Luo et al., 2024a).
+Figure 2.
+Overall illustration of process annotation pipeline of GraphSilo (a, b) and the strategies of utilizing GraphPRM to strengthen LLMs on graph reasoning, including scaling inference-time computation (c) and reinforcement learning (d).
 2.3. Task Formulation
@@
 It is also the only dataset that includes step-level supervision, offering a valuable resource for graph reasoning research.
-The detailed comparison with previous related benchmarks is shown in Table 2.
+We report detailed dataset statistics in Table 9.
 3.2. Supervised Training of GraphPRM
@@
   • RQ6: What is the influence of parameter sizes on the interplay between policy model and GraphPRM?
+More results and analysis about GraphPRM on different models and graph densities are presented at Appendix B.
 4.1. Experimental Settings
@@
 In DPO training, the learning rate is 5e-6, and the preference hyperparameter $\beta$ is set to 0.1.
-Figure 4.
-Best-of-N Performance of GraphPRM on four math datasets with ranging generation budget (number of solution candidates).
-The policy model is Qwen2.5-Math-1.5B-Instruct and the answer aggregation strategy is PRM-Last.
 4.2. Performance of GraphPRM (RQ1)
@@
 Besides, GraphPRM achieves a general better performance transferring to those out-of-domain tasks (Neighbor, BFS, Cycle) that are not included in the training dataset of GraphPRM.
-Dataset Transfer In addition to our GraphSilo, we also evaluate GraphPRM on the widely used GraphWiz dataset.
-We select five most representative tasks (Connect, Topology, Shortest, Hamilton, and Subgraph) across three different difficulty levels and directly test GraphPRM without using example prompts or additional training.
-As shown in Table 4, GraphPRM demonstrates significant improvements in LLMs through inference-time computation, highlighting its effectiveness across datasets with diverse distributions.
+Comparison against ORMs/Math-PRMs As GraphPRM is the first PRM specific for GCP tasks, to further demonstrate the effectiveness of GraphPRM compared to ORM, we train an additional G-ORM on GraphSilo without step-wise labels for GCP tasks (only use outcome reward for supervision).
+As reported in Table 4, we can observe that though G-ORM outperforms SC, it lags behind GraphPRM on both in-domain and out-domain graph reasoning tasks, which further prove the effectiveness of GraphPRM.
+Beyond this, we also investigate the performance of representative math-specific PRMs on GraphSilo, and Table 4 shows that math-specific PRMs cannot easily transfer to graph reasoning tasks.
+We conclude this phenomenon for the following reasons: 1) Graph reasoning tasks involve diverse reasoning patterns (e.g., topological, logical, and computational reasoning) that inherently require an understanding of graph structures, which is less pronounced in mathematical problem-solving.
+2) The difficulty of graph problems often scales exponentially with the size and complexity of the graph.
+Methods In-domain Out-domain Average
+Self-Consistency 37.03 56.67 41.56
+G-ORM 37.40 56.33 41.77
+G-ORM + Self-Consistency 37.77 57.00 42.21
+Math-Shepherd-7B (Wang et al., 2024b) 36.23 54.89 40.54
+Math-PSA-7B (Wang et al., 2024a) 37.30 56.44 41.72
+GraphPRM (Best-of-N) 38.30 58.44 42.95
+GraphPRM (Beam Search) 41.57 57.67 45.28
+Table 4.
+Comparison among GraphPRM, G-ORM and math-specific PRMs on GraphSilo.
+The policy model is Qwen2.5-7B-Instruct.
+The generation budget during verification is set to 8.
+The best results are in Purple.
 Verification Strategy Easy Medium Hard Average
@@
 Table 5. Performances (%) of GraphPRM on GraphWiz dataset in five representative tasks. The generation budget during verification is set to 8. The best results are in Purple.
+Dataset Transfer In addition to our GraphSilo, we also evaluate GraphPRM on the widely used GraphWiz dataset.
+We select five most representative tasks (Connect, Topology, Shortest, Hamilton, and Subgraph) across three different difficulty levels and directly test GraphPRM without using example prompts or additional training.
+As shown in Table 5, GraphPRM demonstrates significant improvements in LLMs through inference-time computation, highlighting its effectiveness across datasets with diverse distributions.
 Generation Budget We further evaluate GraphPRM on LLMs with varying numbers of generation samples to analyze the impact of the generation budget (representing computational effort or token usage per question) on inference-time scaling.
@@
 4.3. Transferability on Math Problems (RQ2)
+Figure 5.
+Best-of-N Performance of GraphPRM on four math datasets with ranging generation budget (number of solution candidates).
+The policy model is Qwen2.5-Math-1.5B-Instruct and the answer aggregation strategy is PRM-Last.
 To investigate the zero-shot transferability of GraphPRM on mathematical reasoning tasks, we conduct experiments on four classic math datasets.
@@
 This dynamic adaptability is equally important when dealing with open-ended and variable mathematical problems.
+Comparison against Math-PRMs on math datasets We further compare GraphPRM with two representative math-specific PRMs: Math-Shepherd-Mistral-7B (Wang et al., 2024b) and MATH-PSA-7B (Wang et al., 2024a).
+Results reported in Table 6 shown that GraphPRM yield strong and competitive performance on math reasoning benchmarks.
+With the observation in Table 4 that math-specific PRMs cannot easily transfer to graph reasoning tasks, it further demonstrates the uniqueness and effective transferability of GraphPRM.
+Methods GSM8K MATH
+Best-of-N Beam Search Best-of-N Beam Search
+Self-Consistency 89.08 89.08 75.40 75.40
+Math-Shepherd-7B 90.09 89.08 75.90 76.00
+Math-PSA-7B 90.50 90.25 76.30 75.40
+GraphPRM 89.92 90.54 76.00 76.30
+Table 6.
+Comparison among GraphPRM and math-specific PRMs on math datasets.
+The policy model is Qwen2.5-Math-1.5B-Instruct.
+The generation budget during verification is set to 8.
+The best results are in Purple.
 4.4. Performance of Post-trained LLMs with Process Supervision (RQ3)
@@
 It aligns well with the static, fixed-pool nature of Best-of-N; 2) Monte-Carlo labels excel in the Beam Search strategy because their diversity and variability allow the iterative refinement process to explore a wider solution space, maximizing the potential for improvement; 3) Combining both comprehensive reasoning processes with fine-grained step-level labels from these two approaches further enhance the effectiveness of GraphPRM.
-Methods Best-of-N Beam Search
-Qwen2.5-7B-Instruct (Self Consistency) 37.03 37.03
-+ GraphPRM 38.30 41.57
-+ GraphPRM - Task-oriented Labels 37.96 36.03
-+ GraphPRM - Monte-carlo Labels 37.36 39.70
-Table 6. Performances (%) of GraphPRM trained on different set of step-wise process labels. The verification is based on 8 candidates. The best results are in Purple.
 4.7. Influence of Pre-trained Model Sizes (RQ6)
@@
 These results substantiate that we should utilize a more potent GraphPRM for validating or supervising the policy model.
+Methods Best-of-N Beam Search
+Qwen2.5-7B-Instruct (Self Consistency) 37.03 37.03
++ GraphPRM 38.30 41.57
++ GraphPRM - Task-oriented Labels 37.96 36.03
++ GraphPRM - Monte-carlo Labels 37.36 39.70
+Table 8. Performances (%) of GraphPRM trained on different set of step-wise process labels. The verification is based on 8 candidates. The best results are in Purple.
 4.8. Case Study
@@
 These results underscore the potential of PRMs to bridge reasoning domains and highlight their promise in advancing step-wise reasoning, laying the groundwork for applying graph-based reasoning to broader challenges and advancing LLM capabilities.
+6. Acknowledgements
+We would like to thank all the anonymous reviewers and area chairs for their insightful and valuable comments.
+We also thank the support of the 12th Baidu Scholarship.
+This research is supported by National Key Research and Development Program of China Grant No.2023YFF0725100 and Guangdong S&T Program C019.
 Appendix A More Dataset Details
@@
 Appendix B Additional Experimental Results
-In this section, to further analysis our proposed GraphPRM, we evaluate GraphPRM by addressing the following question: ”How does GraphPRM perform on other GCP-specific backbones?”
-Performance of GraphPRM on GraphWiz model In addition to our proposed GraphSilo, we conduct experiments on GraphWiz dataset with Qwen2.5-7B-Instruct, Qwen2.5-Math-7B-Instruct and LLaMA-3.1-8B-Instruct in Section 4.2.
+In this section, to further analysis our proposed GraphPRM, we evaluate GraphPRM by addressing the following extra research questions:
+  • RQ7: How does GraphPRM perform on other GCP-specific backbones?
+  • RQ8: How does graph density affect the performance of GraphPRM on LLMs during inference time?
+B.1. Performance of GraphPRM on GraphWiz model (RQ7)
+In addition to our proposed GraphSilo, we conduct experiments on GraphWiz dataset with Qwen2.5-7B-Instruct, Qwen2.5-Math-7B-Instruct and LLaMA-3.1-8B-Instruct in Section 4.2.
 Besides above, we further evaluate the transferbility of GraphPRM on other GCP-specific model like GraphWiz.
@@
 Figure 10. Best-of-N Performance of GraphPRM on GraphWiz with different backbones. We adopt GraphWiz-DPO trained on LLaMA2-7B. Verification is based on 8 candidates.
+B.2. Impact of Graph Density (RQ8)
+Graph density can be a crucial factor influencing the performance of GraphPRM on graph reasoning tasks.
+We regard Edge Density $\left(x=\frac{2|E|}{n(n-1)}\right)$ as the reflection of graph density, and we construct new test set with three-level density: Low ($0<x\leq 0.33$), Middle ($0.33<x\leq 0.67$) and High ($0.67<x\leq 1$).
+Results with different graph densities are shown in figure 11.
+It can be observed that accuracy decreases as graph density increases, but the model with GraphPRM consistently outperforms that with Self-Consistency, demonstrating its robustness across graphs with different densities.
+Besides, as graph density increases, the performance gained via GraphPRM (Best-of-8) decreases.
+We conclude this is the reason that: Graphs with higher density contain more invalid edges or noisy edges, bringing more challenges for GraphPRM to identify essential steps.
+This further highlights the importance of employing GraphPRM to supervise step-wise graph reasoning.
+Figure 11. Best-of-N Performance of GraphPRM on GraphSilo with different graph densities. The policy model is Qwen2.5-7B-Instruct and the verification is based on 8 candidates.
 Appendix C More Cases